# Exploring hidden priors when interpreting gravitational wave and electromagnetic probes of the nuclear equation of state

### v2

@@ IV Conclusions
In this proof-of-concept study, we have investigated how several often-obscured sources of prior information propagate into statements about the nuclear equation of state. Most notably, we have demonstrated that hidden astrophysical population assumptions about the source responsible for GW170817/AT2017gfo directly propagate into substantially different conclusions, both about the radius of a fiducial neutron star and more broadly about the nuclear equation of state. In doing so, we have clarified how EoS inference requires and explicitly depends on strong assumptions about each pertinent neutron star population. Using two EoS parameterizations, we show how severely conclusions about the nuclear EoS can depend on parameterizations and prior choices. To concretely illuminate the impact of the prior, our hyperparameter plots indicate marginal EoS likelihoods; in every analysis, many EoSs with the highest marginal likelihood are not within our hyperparameter posterior, being disfavored by the hyperparameter prior (i.e. the priors on the two spectral parameterizations). Using the marginal likelihood to separate the impact of EoS priors, we show many EoS in our two model families are both consistent with current pulsar and GW observations but support very large NS masses. Accounting for these hidden systematics, we argue the nuclear equation of state is much less well constrained by extant observations than previously reported, particularly by investigations which stack multiple heterogeneous measurements whose relative systematics are not well understood.
@@
Our conclusions are compatible with other recent reassessments of EoS-inference-adjacent observational and theoretical systematics [162]. Notably, recent re-examination of NICER observations of J0030+4051 [163] unveils a wider range of possible interpretations than the widely-used unimodal likelihood employed as input in this work. Our conclusions are also compatible with a recent galactic pulsar observations of J0437, as summarized in an approximate analysis in Appendix E [142].
@@
We have methodically investigated the impact on EoS inference by several external sources of information: galactic pulsar measurements of mass, mass and radius, including the HESS measurement of PSR J1731, GW measurements from GW170817, the nuclear symmetry energy implied by PREX, an upper bound on the NS maximum mass, motivated by energetics of AT2017gfo, and discussed proposed lower bounds on $\tilde{\Lambda}$ also motivated by this event. Using the (range of the) marginal log-likelihoods involved as a diagnostic of information provided by each measurement we argue that the primary measurements (galactic pulsars, including the HESS source; GW measurements; and the symmetry energy) are remarkably consistent for our model family, particularly when allowing for systematic uncertainty in the interpretation of GW170817 due to hidden astrophysical population assumptions. This consistency suggests observations presently only coarsely constrain the EoS space, leaving limited opportunity for more flexible EoS to arrive at qualitatively different conclusions. Thus, rather than use modest tension between different observations (e.g., fidicual GW versus NICER) as a mechanism to draw narrower conclusions about the NS radius, which requires well-controlled systematics, we instead argue these tensions point to wider posterior conclusions than previously reported. Specifically, taking the union of our credible intervals between the most optimistic and pessimistic scenarios allowing for both choices for EoS family, we argue $R_{1.4}$ is between $11~{}{\rm km}$ and $14~{}{\rm km}$; see Figures 10 and 6. This range is qualitatively consistent with a simple, pessimistic heuristic interpretation of how systematics limit our observations, when analyzed jointly and accounting for systematics: individual measurements’ maximum likelihood estimates bracket the available range. We emphasize that these optimistic and pessimistic conclusions are driven in no small part by assumptions: which events and interpretations are used, with what EoS family and priors.
In our proof-of-concept study, we have intentionally adopted a fixed low-density EoS and thus a largely known NS radius at low mass. By adopting such a strong prior, we guarantee that low-mass NS will have the expected radius and limit the possibility that even extreme choices for EoS parameters can be in tension with observations. In other words, by fixing the low-density EoS, we increase the chance that the radius $R_{1.4}$ will be consistent with most measurements, leaving other phenomenological coordinates like $M_{max}$ as free (and observationally unbounded) parameters. While in this work for simplicity we have adopted a fixed low-dimensional nuclear equation of state, we recognize that the high precision of future EoS constraints requires a more flexible, powerful framework that incorporates more dimensions and allows for more flexible physics like phase transitions.
To carry out our calculations, we have deployed an updated version of the RIFT pipeline capable of efficiently performing hierarchical inference with multiple objects, multiple messengers, and even flexible source populations to infer a joint equation of state. We built our analysis pipeline using fixed hierarchical inputs: here, approximated NICER likelihoods and high-precision marginal GW likelihoods. In future work we will generalize this pipeline to reassess the hierarchical inputs and request further refinement when needed (e.g., more marginal likelihood evaluations from the GW analysis, to better resolve the low-$\tilde{\Lambda}$ region if needed).
@@
In Appendix B, we outline a simple estimate of how well future GW surveys can constrain the nuclear equation of state. This estimate demonstrates that most individual NS observations will directly inform the nuclear EoS, not just the loudest few. This estimate also demonstrates that, due to the strong trend of tidal deformability with radius, the least massive NS should dominate conclusions about the EoS, excepting only that nature produces them too infrequently. Additionally, this appendix illustrates how EOS constraints will depend on two key assumptions held fixed in the main text: the assumption of a fixed, uniform NS mass distribution on the one hand, and on the number and distance distribution on the other.
The most surprising new result in our study is the strong dependence of GW-inferred tidal deformability on astrophysical prior assumptions. In addition to a fiducial analysis of GW170817, we also analyze this event using two well-motivated binary NS population models. In one model, we require GW170817 be more similar to the expected binary neutron star population, based on theoretical models for their formation: a comparable mass ratio, and with spins preferentially positively aligned with the orbital angular momentum. In the other model, designed to produce (much) larger amounts of radioactive ejecta, we require a highly asymmetric binary (potentially rapidly spinning). These two prior assumptions notably change the conclusions derived from GW170817 and hence the conclusions derived when combining this with other events. Though the latter of our two scenarios seems in tension with the long-lived galactic BNS population, we note that a rapidly-merging subpopulation could have different population properties, and offers an attractive resolution to the long-term challenge of forming r-process elements. Conversely, if we adopt a conventional galactic BNS origin for 170817 and incorporate all available measurements, we conclude $R_{1.4}\simeq 11.8~{}{\rm km}-12.9~{}{\rm km}$.
@@
For simplicity, in this analysis we have omitted several pertinent areas of investigations. We have notably adopted a single GW waveform model, eliding the effect of waveform modeling systematics; see Appendix C for an analysis using TEOBResumS showing that changing the waveform used for EOS inference has a small effect on baseline conclusions, in contrast to changes in the prior assumptions used. Of course, GW waveform models are only calibrated over a certain range of mass ratio, spins, and tides; for the purposes of illustration, at least one of our calculations (the use of high-mass-ratio binaries) stretches the regime of validity of the NRTidalV2 approximation. That said, our headline result – that NS population priors can strongly impact EOS inference, by changing the interpretation of GW170817 – also follows from another example, using comparable masses and positive spins, well within the calibration regime of the NRTidal method. We therefore anticipate the effect size of NS population priors to remain large even as GW waveform models improve. Also, given substantial modeling systematics, we have also only adopted toy models for the information provided from kilonova observations which limit the possible BNS mass, spin, and EoS. We have taken all existing fiducial galactic pulsar measurements at face value; reanalyses of these observations may point to larger systematics and broader posteriors than applied here [163]. Finally, we have adopted an intentionally simple EoS model family and prior, omitting often-discussed constraints available at low density (e.g., from chiral effective field theory and from state-of-the-art crust models) and at high density (e.g, that high-density EoS should asymptotically attain the perturbative-QCD limit of $c_{s}\rightarrow 1/\sqrt{3}$ approaching from below [164, 165]). Our choice of these simplifications enables us to focus exclusively on the often strong and hidden impact that hidden astrophysical priors produce.
@@
The multifactor iterative inference algorithm employed in this work is available as the HyperPipe component of the RIFT python software package, available at https://pypi.org/project/RIFT/. Instructions for this approach are available at [166]. Appendix D illustrates how the hyperpipeline accumulaties information iteratively, in a manner similar to RIFT.
@@ Acknowledgements.
We thank Jocelyn Read for constructive feedback during the development of this work. A.K. thanks Wolfgang Kastaun for help with using RePrimAnd. We thank the NICER collaboration, in particular M. Coleman Miller, for providing us with data for certain figures. A.K. acknowledges support from National Science Foundation (NSF) Grant No. AST-1909534. ROS acknowledges support from NSF Grant No. AST-1909534, NSF Grant No. PHY-2012057, and the Simons Foundation. A.B.Y. acknowledges support from NSF Grant No. PHY-2012057. The EoS parameterizations were developed under my NSF Grant No. PHYS-2011874. We are grateful for the computational resources provided by the LIGO Laboratory, supported via National Science Foundation Grants PHY-0757058 and PHY-0823459. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. The code settings and data products needed to reproduce and validate our results are available on request.
@@ Appendix A NICER data and Gaussian approximation
The NICER data releases [167, 168, 169, 170] provide weighted posterior samples $(x_{k},w_{k})$ derived from comparing their pulse models and neutron star mass-radius combinations with observed X-ray photons. We fit a conventional two-dimensional normal distribution to these weighted samples. For J0030, which lacks any hard edges introduced by causality constraints in the $(M,R)$ plane, we use the weighted empirical mean $\bar{x}=\sum_{k}w_{k}x_{k}/\sum_{k}w_{k}$ and weighted empirical covariance matrix to infer the parameters of the best-fitting gaussian. More generally, however, we optimize the mean $\mu$ and covariance matrix $\Sigma$ of a standard normal distribution, according to
$\ell(\mu,\Sigma)=\sum_{k}w_{k}\ln p(x_{k}|\mu,\Sigma)~{}~{}.$ (13)
In the limit of many weighted samples constructed by importance sampling, with $w_{k}=L(x_{k})$, the quantity $\ell$ will approach the KL divergence between the (truncated) Gaussian and the full posterior. As a check on our results and their stability, we repeat our analysis with fair draws from the weighted samples, using the expression above with $w_{k}=1$. The best fitting normal distribution in $(R/{\rm km},~{}M/{\rm M_{\odot}})$ has a mean $(13.34537396,~{}1.44968231)$ and a covariance matrix
$\begin{bmatrix}1.52687831&0.18343073\\ 0.18343073&0.02778826\end{bmatrix}~{}~{}~{}.$
@@ Appendix B Understanding the impact of many observations
Future instruments will enable very high precision constraints on the nuclear equation of state, through high precision constraints on $\Lambda(m)$. Starting with the pioneering work using the Fisher matrix [171], several groups have investigated how well EoS or radius constraints will improve as observations accumulate, using various levels of approximation; see, e.g., [104, 24, 172, 173, 174] and references therein. Unfortunately, many misconceptions about these constraints persist in the literature. Notably, most groups assert that the loudest few observations dominate the overall constraint. As discussed in Wysocki et al. [93] however, not only do the many weak observations make critical contributions to the overall EoS constraint, but also we must use these weak observations to help simultaneously reconstruct the NS population.
Motivated by the need for practical projections for third-generation instruments and the many-detection era, in this appendix we provide a simple ansatz allowing simple estimates showing how statistical measurement errors on the nuclear equation of state will improve as more observations accumulate. Our argument has five key ingredients. First, we will assume all binaries are equal mass. Second, following Ref. [136] we will assume $\Lambda(m)=\Lambda_{*}(m/m_{*})^{-p}$ with $p=6$, where $\Lambda_{*}$ (the equation of state parameter to be measured with many observations) is the value of the tidal deformability at a reference mass $m_{*}$ (assumed to be at or close to $m_{*}=1.4~{}{\rm M_{\odot}}$). Third, following the discussion in Section III.A of Ref. [93], we will argue the marginal one-dimensional measurement accuracy for the tidal deformability scales as $\sigma_{\Lambda}=\sigma_{*}(m/m_{*})^{-1}$ versus binary mass, adopting a fiducial uncertainty for the faintest source used in our analysis of $\sigma_{*}\simeq 200$; see also, e.g., Fig. 1 in [172]. Fourth, we will assume that each independent observation’s marginal tidal deformability has a log-likelihood of the form
$\Gamma_{**,1}=\Big{(}\frac{\rho}{8}\Big{)}^{2}\sigma_{\lambda}^{-2}\Big{(}\frac{d\Lambda}{d\Lambda_{*}}\Big{)}^{2}=\Big{(}\frac{\rho}{8}\Big{)}^{2}\sigma_{*}^{-2}\Big{(}\frac{m}{m_{*}}\Big{)}^{2-2p}~{}~{}~{}.$ (14)
@@
where in the second line we have replaced the discrete sum by an ensemble average ($\sum_{k}\cdot=N{\left\langle\cdot\right\rangle}$) and for convenience in our ansatz we assume the angular, distance, and mass distributions of the $N$ detected sources are independent. The average ${\left\langle w^{2}\right\rangle}$ of the geometrical factors depends on the assumed network beampattern (and, for realistic heterogeneous networks, their duty cycle), but is well known to be ${\left\langle w^{2}\right\rangle}=2/5$ for the single-interferometer scenario; see Refs. [175, 176, 177] for discussion. Unless otherwise noted, for the purposes of discussion we will assume the single-interferometer beampattern factor, a uniform NS mass distribution between $1~{}{\rm M_{\odot}}$ and $2~{}{\rm M_{\odot}}$, and use ${\left\langle\Big{(}\frac{d_{L,*}}{d_{L}}\Big{)}^{2}\right\rangle}\simeq 1$ as appropriate to a third-generation network. A larger value of this expression corresponds to smaller marginal measurement accuracy $\sigma_{\Lambda_{*,N}}\equiv 1/\sqrt{\Gamma_{**,N}}$ on the fiducial tidal deformability $\Lambda_{*}$.
@@
Our expression provides an even stronger conclusion about the relative importance of different parts of the mass distribution: the lowest mass binaries overwhelmingly dominate the mass average ${\left\langle m^{11/3-2p}\right\rangle}\simeq{\left\langle m^{-8}\right\rangle}$ (more precisely = $\frac{1-2^{-(8-2/3)}}{9-2/3}$ for the average for a uniform mass distribution from $1~{}{\rm M_{\odot}}$ to $2~{}{\rm M_{\odot}}$). In retrospect, the lowest mass binaries must dominate: $\Lambda(m)$ is vastly larger at the lowest allowed masses. If any binaries exist at these masses their observations will dominate our constraints on the equation of state because their large tidal effects will be easily differentiated from the noise.
@@
Our simple estimate is comparable to (but calibrated above to be slightly more optimistic than) detailed calculations produced in other studies [178, 104, 172]; for example, ref. [104] argues the 90% credible interval is $\Delta\Lambda_{*}\simeq 910/\sqrt{N}$ and thus that the comparable $68\%$ credible interval would be $520/\sqrt{N}$. The aforementioned measurement accuracy on $\Lambda_{*}$ implies a relative measurement accuracy on $\ln R_{*}$ of $\Lambda_{*}^{-1}|d\ln R_{*}/d\ln\Lambda_{*}|/\sqrt{\Gamma_{**,N}}$. To evaluate this expression and put the result in context, we’ll assume for the purposes of illustration the MPA1 equation of state, which produces for $m_{*}=1.4~{}{\rm M_{\odot}}$ the fiducial pair $(\Lambda_{*},R_{*}/{\rm km})=(513,12.5)$. For simplicity again following Zhao and Lattimer [136], we can evaluate this change-of-variable expression by adopting the same scaling $\Lambda=a(m/R)^{-p}$ assumed above for some constant $a\simeq 0.0085$, so $d\ln R_{*}/d\ln\Lambda_{*}=p$. We therefore estimate the measurement accuracy $\sigma_{R_{*},N}$ in $R_{*}$ produced by stacking $N$ measurements to be
@@
which is close to (but slightly wider than) the corresponding estimate presented in [104]. Our simple estimates can be easily updated (and compared to the latest best-fit value $\Lambda_{*}$) as our understanding of the NS mass distribution and measurement accuracy improves. They can likewise be easily updated to include both GW and X-ray observations adding in quadrature.
Our analysis in this appendix suggests important trends in tension with previous preliminary investigations. Many groups over several years have argued that only the loudest handful of observations will dominate overall EoS inference [171, 179]. Our analytic argument suggests a comparable and potentially even dominant contribution from more numerous fainter events.
In the discussion in this appendix, we have only provided estimates for the statistical accuracy $\sigma_{\Lambda_{*},N}$ and $\sigma_{R_{*},N}$ for the tidal deformability and radius of a fiducial neutron star, respectively, after combining $N$ events. Given proposed network sensitivity models and reference, these expressions can be immediately translated to the operating time needed to constrain these quantities to any target precision. For example, in terms of the proposed detection rate of useful events (scaled to $10^{5}/{\rm yr}$, assuming all events contribute) we find for the simple estimates discussed here that a target accuracy of $\Delta R_{*}=10~{}{\rm m}$ requires an observing time scaling with event rate and fiducial neutron star radius as

### v3

@@ 5 Conclusions
In this study we have investigated how implicit assumptions propagate into statements about the nuclear equation of state. In particular, we have demonstrated that astrophysical population assumptions about the source responsible for GW170817/AT2017gfo directly propagate into substantially different conclusions, both about the radius of a fiducial neutron star and consequently the nuclear equation of state. In doing so, we have shown how EoS inference requires and explicitly depends on strong assumptions about each pertinent neutron star population. The impact of astrophysical population assumptions is of shifting the $R_{1.4\penalty 10000\ {\rm M_{\odot}}}$ by $\pm 0.5\penalty 10000\ {\rm km}$, whereas individual observations only shift $R_{1.4\penalty 10000\ {\rm M_{\odot}}}$ by $\pm 0.1\penalty 10000\ {\rm km}$. Using two EoS parameterizations, we show how severely conclusions about the nuclear EoS can depend on parameterization prior choices. To illuminate the impact of the prior, the hyperparameter plots indicate marginal EoS likelihoods; in every analysis, many EoSs with the highest marginal likelihood are not within the hyperparameter posterior, being disfavored by the hyperparameter prior (i.e. the priors on the two spectral parameterizations). Using the marginal likelihood to separate the impact of EoS priors, we show many EoS in our two model families are both consistent with current pulsar and GW observations but support very large NS masses. Accounting for these systematics, we argue that for the first few observational constraints, which previously seemed to establish a consensus about the EoS, the nuclear equation of state is much less well constrained than previously reported, particularly by investigations which stack multiple heterogeneous measurements whose relative systematics and particular population assumptions are not well understood.
The most surprising new result in our study is the strong dependence of GW-inferred tidal deformability on astrophysical prior assumptions. In addition to a fiducial analysis of GW170817, we also analyze this event using two well-motivated binary NS population models. In one model, we require GW170817 be more similar to the expected binary neutron star population, based on theoretical models for their formation: a comparable mass ratio, and with spins preferentially positively aligned with the orbital angular momentum. In the other model, designed to produce (much) larger amounts of radioactive ejecta, we require a highly asymmetric binary (potentially rapidly spinning). These two prior assumptions notably change the conclusions derived from GW170817 and hence the conclusions derived when combining this with other events. Though the latter of our two scenarios seems in tension with the long-lived galactic BNS population, we note that a rapidly-merging subpopulation could have different population properties, and offers an attractive resolution to the long-term challenge of forming r-process elements. Conversely, if we adopt a conventional galactic BNS origin for 170817 and incorporate all early-era measurements reviewed in this paper, we conclude $R_{1.4}\simeq 11.8\penalty 10000\ {\rm km}-12.9\penalty 10000\ {\rm km}$.
@@
Our conclusions are compatible with other recent reassessments of EoS-inference-adjacent observational and theoretical systematics (2023ApJ...944....7S). Notably, recent re-examination of NICER observations of J0030+4051 (2024ApJ...961...62V) unveils a wider range of possible interpretations than the widely-used unimodal likelihood employed as input in this work. Our conclusions are also compatible with a recent galactic pulsar observations of J0437, as summarized in an approximate analysis in Appendix E (2024ApJ...971L..20C).
@@
We have methodically investigated the impact on EoS inference by several external sources of information: galactic pulsar measurements of mass, mass and radius, including the HESS measurement of PSR J1731, GW measurements from GW170817, the nuclear symmetry energy, an upper bound on the NS maximum mass motivated by energetics of AT2017gfo, and discussed proposed lower bounds on $\tilde{\Lambda}$ also motivated by this event. Using the (range of the) marginal log-likelihoods involved as a diagnostic of information provided by each measurement we argue that the primary measurements (galactic pulsars, including the HESS source; GW measurements; and the symmetry energy) are remarkably consistent for our model family, particularly when allowing for systematic uncertainty in the interpretation of GW170817 due to astrophysical population assumptions. This consistency suggests observations presently only coarsely constrain the EoS space, leaving limited opportunity for more flexible EoS to arrive at qualitatively different conclusions. Thus, rather than use modest tension between different observations (e.g., fidicual GW versus NICER) as a mechanism to draw narrower conclusions about the NS radius, which requires well-controlled systematics, we instead argue these tensions point to wider posterior conclusions than previously reported. Specifically, taking the union of our credible intervals between the most optimistic and pessimistic scenarios allowing for both choices for EoS family, we argue $R_{1.4}$ is between $11\penalty 10000\ {\rm km}$ and $14\penalty 10000\ {\rm km}$; see Figures 10 and 6. This range is qualitatively consistent with a simple, pessimistic, heuristic interpretation of how systematics limit our observations, when analyzed jointly and accounting for systematics: individual measurements’ maximum likelihood estimates bracket the available range. Preliminary interpretations of J0614 are qualitatively consistent with our pessimistic conclusion. We emphasize that these optimistic and pessimistic conclusions are driven in no small part by assumptions: which events and interpretations are used, with what EoS family and priors.
@@
In Appendix B, we outline a simple estimate of how well future GW surveys can constrain the nuclear equation of state. This estimate demonstrates that most individual NS observations will directly inform the nuclear EoS, not just the loudest few. This estimate also demonstrates that, due to the strong trend of tidal deformability with radius, the least massive NS should dominate conclusions about the EoS, excepting only that nature produces them too infrequently. Additionally, the appendix illustrates how EOS constraints will depend on two key assumptions held fixed in the main text: the assumption of a fixed, uniform NS mass distribution and the number and distance distribution.
@@
For simplicity, in this analysis we have omitted several pertinent areas of investigations. We have notably adopted a single GW waveform model, eliding the effect of waveform modeling systematics; see Appendix C for an analysis using TEOBResumS showing that changing the waveform used for EOS inference has a small effect on baseline conclusions, in contrast to changes in the prior assumptions used. Of course, GW waveform models are only calibrated over a certain range of mass ratio, spins, and tides; for the purposes of illustration, at least one of our calculations (the use of high-mass-ratio binaries) stretches the regime of validity of the NRTidalV2 approximation. That said, our headline result – that NS population priors can strongly impact EOS inference, by changing the interpretation of GW170817 – also follows from another example, using comparable masses and positive spins, well within the calibration regime of the NRTidal method. We therefore anticipate the effect size of NS population priors to remain large even as GW waveform models improve. Also, given substantial modeling systematics, we have also only adopted toy models for the information provided from kilonova observations which limit the possible BNS mass, spin, and EoS. We have taken all existing fiducial galactic pulsar measurements at face value; reanalyses of these observations may point to larger systematics and broader posteriors than applied here (2024ApJ...961...62V). Finally, we have adopted an intentionally simple EoS model family and prior, omitting often-discussed constraints available at low density (e.g., from chiral effective field theory and from state-of-the-art crust models) and at high density (e.g, that high-density EoS should asymptotically attain the perturbative-QCD limit of $c_{s}\rightarrow 1/\sqrt{3}$ approaching from below (2010PhRvD..81j5021K; 2020NatPh..16..907A)). Our choice of these simplifications enables us to focus exclusively on the often strong and impact that astrophysical priors produce.
@@
The multifactor iterative inference algorithm employed in this work is available as the HyperPipe component of the RIFT python software package, available at https://pypi.org/project/RIFT/. Instructions for this approach are available at (rift-docs). Appendix D illustrates how the hyperpipeline accumulates information iteratively in a manner similar to RIFT.
@@ Acknowledgements.
We thank Jocelyn Read for constructive feedback during the development of this work. A.K. thanks Wolfgang Kastaun for help with using RePrimAnd. We thank the NICER collaboration, in particular M. Coleman Miller, for providing us with data for certain figures. A.K. acknowledges support from National Science Foundation (NSF) Grant No. AST-1909534. ROS acknowledges support from NSF Grant No. AST-1909534, NSF Grant No. PHY-2012057, and the Simons Foundation. A.B.Y. acknowledges support from NSF Grant No. PHY-2012057. The EoS parameterizations were developed under my NSF Grant No. PHYS-2011874. We are grateful for the computational resources provided by the LIGO Laboratory, supported via National Science Foundation Grants PHY-0757058 and PHY-0823459. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. The code settings needed to reproduce and validate our results are available on https://pypi.org/project/RIFT/ and https://github.com/oshaughn/research-projects-RIT/.
@@ Appendix A NICER data and Gaussian approximation
The NICER data releases (riley_thomas_e_2019_3386449; miller_m_c_2019_3473466; riley_thomas_e_2021_4697625; miller_m_c_2021_4670689) provide weighted posterior samples $(x_{k},w_{k})$ derived from comparing their pulse models and neutron star mass-radius combinations with observed X-ray photons. We fit a conventional two-dimensional normal distribution to these weighted samples. For J0030, which lacks any hard edges introduced by causality constraints in the $(M,R)$ plane, we use the weighted empirical mean $\bar{x}=\sum_{k}w_{k}x_{k}/\sum_{k}w_{k}$ and weighted empirical covariance matrix to infer the parameters of the best-fitting gaussian. More generally, however, we optimize the mean $\mu$ and covariance matrix $\Sigma$ of a standard normal distribution, according to
$\ell(\mu,\Sigma)=\sum_{k}w_{k}\ln p(x_{k}|\mu,\Sigma)\penalty 10000\ \penalty 10000\ .$ (13)
In the limit of many weighted samples constructed by importance sampling, with $w_{k}=L(x_{k})$, the quantity $\ell$ will approach the KL divergence between the (truncated) Gaussian and the full posterior. As a check on our results and their stability, we repeat our analysis with fair draws from the weighted samples, using the expression above with $w_{k}=1$. The best fitting normal distribution in $(R/{\rm km},\penalty 10000\ M/{\rm M_{\odot}})$ has a mean $(13.34537396,\penalty 10000\ 1.44968231)$ and a covariance matrix
$\begin{bmatrix}1.52687831&0.18343073\\ 0.18343073&0.02778826\end{bmatrix}\penalty 10000\ \penalty 10000\ \penalty 10000\$.
@@ Appendix B Understanding the impact of many observations
Future instruments will enable very high precision constraints on the nuclear equation of state, through high precision constraints on $\Lambda(m)$. Starting with the pioneering work using the Fisher matrix (2015PhRvD..91d3002L), several groups have investigated how well EoS or radius constraints will improve as observations accumulate, using various levels of approximation; see, e.g., 2020PhRvD.101l3007L; 2020PhRvR...2c3514P; 2023arXiv230705376H; 2024CQGra..41v5003B; 2023ApJ...955...45F and references therein. Unfortunately, many misconceptions about these constraints persist in the literature. Notably, most groups assert that the loudest few observations dominate the overall constraint. As discussed in 2020arXiv200101747W however, not only do the many weak observations make critical contributions to the overall EoS constraint, but also we must use these weak observations to help simultaneously reconstruct the NS population.
Motivated by the need for practical projections for third-generation instruments and the many-detection era, in this appendix we provide a simple ansatz allowing simple estimates showing how statistical measurement errors on the nuclear equation of state will improve as more observations accumulate. Our argument has five key ingredients. First, we will assume all binaries are equal mass. Second, following Zhao:2018nyf we will assume $\Lambda(m)=\Lambda_{*}(m/m_{*})^{-p}$ with $p=6$, where $\Lambda_{*}$ (the equation of state parameter to be measured with many observations) is the value of the tidal deformability at a reference mass $m_{*}$ (assumed to be at or close to $m_{*}=1.4\penalty 10000\ {\rm M_{\odot}}$). Third, following the discussion in Section III.A of (2020arXiv200101747W), we will argue the marginal one-dimensional measurement accuracy for the tidal deformability scales as $\sigma_{\Lambda}=\sigma_{*}(m/m_{*})^{-1}$ versus binary mass, adopting a fiducial uncertainty for the faintest source used in our analysis of $\sigma_{*}\simeq 200$; see also, e.g., Fig. 1 in 2023arXiv230705376H. Fourth, we will assume that each independent observation’s marginal tidal deformability has a log-likelihood of the form
$\Gamma_{**,1}=\Big(\frac{\rho}{8}\Big)^{2}\sigma_{\lambda}^{-2}\Big(\frac{d\Lambda}{d\Lambda_{*}}\Big)^{2}=\Big(\frac{\rho}{8}\Big)^{2}\sigma_{*}^{-2}\Big(\frac{m}{m_{*}}\Big)^{2-2p}\penalty 10000\ \penalty 10000\ \penalty 10000\ .$ (14)
@@
where in the second line we have replaced the discrete sum by an ensemble average ($\sum_{k}\cdot=N{\left\langle\cdot\right\rangle}$) and for convenience in our ansatz we assume the angular, distance, and mass distributions of the $N$ detected sources are independent. The average $\left\langle w^{2}\right\rangle$} of the geometrical factors depends on the assumed network beampattern (and, for realistic heterogeneous networks, their duty cycle), but is well known to be ${\left\langle w^{2}\right\rangle}=2/5$ for the single-interferometer scenario; see 1993PhRvD..47.2198F; 1994PhRvD..49.2658C; 2010PhRvD..82j4006O for discussion. Unless otherwise noted, for the purposes of discussion we will assume the single-interferometer beampattern factor, a uniform NS mass distribution between $1\penalty 10000\ {\rm M_{\odot}}$ and $2\penalty 10000\ {\rm M_{\odot}}$, and use ${\left\langle\Big(\frac{d_{L,*}}{d_{L}}\Big)^{2}\right\rangle}\simeq 1$ as appropriate to a third-generation network. A larger value of this expression corresponds to smaller marginal measurement accuracy $\sigma_{\Lambda_{*,N}}\equiv 1/\sqrt{\Gamma_{**,N}}$ on the fiducial tidal deformability $\Lambda_{*}$.
@@
Our expression provides an even stronger conclusion about the relative importance of different parts of the mass distribution: the lowest mass binaries overwhelmingly dominate the mass average ${\left\langle m^{11/3-2p}\right\rangle}\simeq{\left\langle m^{-8}\right\rangle}$ (more precisely = $\frac{1-2^{-(8-2/3)}}{9-2/3}$ for the average for a uniform mass distribution from $1\penalty 10000\ {\rm M_{\odot}}$ to $2\penalty 10000\ {\rm M_{\odot}}$). In retrospect, the lowest mass binaries must dominate: $\Lambda(m)$ is vastly larger at the lowest allowed masses. If any binaries exist at these masses their observations will dominate our constraints on the equation of state because their large tidal effects will be easily differentiated from the noise.
@@
Our simple estimate for the expected statistical uncertainty in this parameter after $N$ observations is comparable to (but calibrated above to be slightly more optimistic than) detailed calculations produced in other studies (2023PhRvC.108b5811R; 2020PhRvD.101l3007L; 2023arXiv230705376H); for example, 2020PhRvD.101l3007L argues the 90% credible interval is $\Delta\Lambda_{*}\simeq 910/\sqrt{N}$ and thus that the comparable $68\%$ credible interval would be $520/\sqrt{N}$. The aforementioned measurement accuracy on $\Lambda_{*}$ implies a relative measurement accuracy on $\ln R_{*}$ of $\Lambda_{*}^{-1}|d\ln R_{*}/d\ln\Lambda_{*}|/\sqrt{\Gamma_{**,N}}$. To evaluate this expression and put the result in context, we’ll assume for the purposes of illustration the MPA1 equation of state, which produces for $m_{*}=1.4\penalty 10000\ {\rm M_{\odot}}$ the fiducial pair $(\Lambda_{*},R_{*}/{\rm km})=(513,12.5)$. For simplicity again following Zhao:2018nyf, we can evaluate this change-of-variable expression by adopting the same scaling $\Lambda=a(m/R)^{-p}$ assumed above for some constant $a\simeq 0.0085$, so $d\ln R_{*}/d\ln\Lambda_{*}=p$. We therefore estimate the measurement accuracy $\sigma_{R_{*},N}$ in $R_{*}$ produced by stacking $N$ measurements to be
@@
which is close to (but slightly wider than) the corresponding estimate presented in (2020PhRvD.101l3007L). Our simple estimates can be easily updated (and compared to the latest best-fit value $\Lambda_{*}$) as our understanding of the NS mass distribution and measurement accuracy improves. They can likewise be easily updated to include both GW and X-ray observations adding in quadrature.
Our analysis in this appendix suggests important trends in tension with previous preliminary investigations. Many groups over several years have argued that only the loudest handful of observations will dominate overall EoS inference (2015PhRvD..91d3002L; PhysRevD.110.043013). Our analytic argument suggests a comparable and potentially even dominant contribution from more numerous fainter events.
In the discussion in this appendix, we have only provided estimates for the statistical accuracy $\sigma_{\Lambda_{*},N}$ and $\sigma_{R_{*},N}$ for the tidal deformability and radius of a fiducial neutron star, respectively, after combining $N$ events. Given proposed network sensitivity models and reference, these expressions can be immediately translated to the operating time needed to constrain these quantities to any target precision. For example, in terms of the proposed detection rate of useful events (scaled to $10^{5}/{\rm yr}$, assuming all events contribute) we find for the simple estimates discussed here that a target accuracy of $\Delta R_{*}=10\penalty 10000\ {\rm m}$ requires an observing time scaling with event rate and fiducial neutron star radius as